# Informed Correctors for Discrete Diffusion Models

### v4

@@ 1 Introduction
Denoising diffusion models are powerful generative models for high-dimensional data (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Ho et al., 2020). The central idea of diffusion models is to gradually corrupt data into noise using a “noising” or “forward” process and then to train a parameterized model (usually a deep neural network) to learn its time reversal, commonly known as the “denoising” or “backward” process. In continuous domains like image generation, the forward process is usually defined by gradual injection of Gaussian noise and scaling that transform the data distribution into a standard normal. The backward process is then approximated by learning the gradient of the log density (also known as the “score function”) of the marginal distribution. To draw samples from a trained model, one first generates Gaussian noise and then simulates the backward process using the score information. Diffusion generative models are currently the dominant framework for image generation, and they can generate high-resolution images with stunning details given prompts from a user (Ramesh et al., 2021; Rombach et al., 2022).
@@
Given the success of diffusion models for continuous data, recent work has explored diffusion modeling in discrete domains (Hoogeboom et al., 2021a; Austin et al., 2021; Campbell et al., 2022; Santos et al., 2023; Zheng et al., 2023; Lou et al., 2023; Shi et al., 2024). Discrete diffusion models can be applied to language (Lou et al., 2023), protein sequences (Watson et al., 2023; Alamdari et al., 2023), graphs (Vignac et al., 2023), and more. Notably, Campbell et al. (2022) developed a general framework for discrete denoising diffusion in continuous time. They formulated the forward and backward processes as continuous time Markov chains (CTMCs), and they learned to predict the distribution over denoised data via an evidence lower bound objective. Concurrent work by Shi et al. (2024); Sahoo et al. (2024) and Ou et al. (2024) focused on discrete diffusion with the absorbing state forward process (a.k.a., masked discrete diffusion), deriving a simple formulation of the ELBO that reduces to a weighted cross-entropy. Shi et al. (2024) also noted that this simple objective provides a unifying perspective to previous works (Campbell et al., 2022; Lou et al., 2023; Benton et al., 2024), all of which are parameterizations and modifications of the same ELBO objective.
@@
Despite conceptual advances in discrete diffusion, the efficiency-accuracy trade-off in simulating the continuous-time backward process still limit their effectiveness. Song et al. (2021) proposed using approximate MCMC correctors to fix the discretization error in simulating the backward process. Campbell et al. (2022) extended the predictor-corrector sampling to discrete diffusion models and discovered that the corrector steps significantly impact generative performance. In practice however, a clear understanding of how corrector steps contribute to sample quality is still missing, and recent works (Lou et al., 2023; Shi et al., 2024; Ou et al., 2024; Sahoo et al., 2024) still use predictor-only samplers for generation.
@@
We show that predictor-corrector schemes can be improved by leveraging different sampling methods. Analyzing the forward-backward corrector used by Campbell et al. (2022), we identify its failure mode in masked diffusion and illustrate it with a simple Markov chain modeling task. To address this shortcoming, we propose the informed corrector — an alternative corrector scheme inspired by adaptive Gibbs sampling that fixes the issues of the forward-backward corrector and explicitly targets low-probability dimensions. Sampling with the informed corrector drastically decreases error rate in our synthetic example, as well as in Text8, where the model makes much fewer spelling errors than baselines. On tokenized ImageNet $256\times 256$, we show that the informed corrector yield diverse, high quality samples competitive with existing diffusion models.
@@ 2.1 Continuous-Time Discrete Diffusion Models
Consider a data vector $x_{0}\in\mathcal{S}^{D}$ sampled from a data distribution $p_{\text{data}}$, where $\mathcal{S}$ is the base space for each component. Denoising diffusion models (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Ho et al., 2020; Song et al., 2021) model this unknown data distribution by progressively adding noise to data samples until their distribution converges to a simple target, such as a uniform distribution. Then, a deep neural network is trained to invert this process and remove the noise. This reverse process forms a generative model from which new samples can be drawn. The noising and generative processes are also called the forward and backward processes, respectively. We will adopt this convention for the rest of the paper.
@@
Campbell et al. (2022) approximated the backward rate by learning a parameterized denoising model $p^{\theta}_{0|t}(x_{0}\mid y)\approx q_{0|t}(x_{0}\mid y)$. Alternatively, Meng et al. (2022) and Lou et al. (2023) observed that the ratio $s_{t}(y)_{x}\triangleq\nicefrac{{q_{t}(x)}}{{q_{t}(y)}}$ plays a similar role as the continuous score function $\nabla\log q_{t}(y)$, up to an additive constant. Therefore, they proposed to learn an approximate score function instead, $s^{\theta}_{t}(y)$.
@@
for a base rate matrix $R_{b}\in\mathbb{R}^{S\times S}$, absorbing the time dependence into the scalar coefficient $\beta(t)$. Here, $x^{\setminus d}$ represents all components of $x$ apart from the $d$-th dimension. Backward rates are also factorized with this forward rate factorization, allowing the objective in ˜2 to be nicely decomposed into a sum over dimensions, as detailed by Campbell et al. (2022).
@@ 2.2 Absorbing Discrete Diffusion
For generic discrete data, two forward processes are most commonly used: the uniform process and the absorbing state process (Austin et al., 2021). We focus on the absorbing state process in this paper, for which the base rate matrix for each dimension can be written as
@@
where we augment the vocabulary set $\mathcal{S}$ by introducing the Mask token, which represents the absorbing state. The absorbing state diffusion is important to consider because it is very commonly used to model sequence data without ordinal structure (e.g., text and protein sequences). It is also intimately connected to masked language models (Austin et al., 2021; Devlin et al., 2018; Hoogeboom et al., 2021b).
@@
Recently, Shi et al. (2024) introduced the MD4 diffusion model, which outperformed previous discrete diffusion models on both text and pixel-level image generation benchmarks. Furthermore, they noted that several continuous time denoising diffusion objectives (including ˜2 ) are mathematically equivalent and can be greatly simplified in the case of masked diffusion. The simplified MD4 loss is:
@@ 2.3 Sampling From the Backward Process
Campbell et al. (2022) proposed to use tau-leaping, an approximate sampling method that allows simultaneous updates to multiple dimensions for a single model evaluation. Tau-leaping assumes that the backward process transition rate is constant between $t-\tau,t)$, regardless of whether states change during that interval. Under this assumption, multiple transition events in the interval can be applied simultaneously. Leverging the structure of masked diffusion, MD4 (Shi et al., 2024) used a simpler approach. Instead of trying to simulate the continuous time process, they directly sampled from the finite-time backward conditional distribution $p^{\theta}_{t-\Delta t\mid t}(\,\cdot\mid x_{t})$. They call this approach ancestral sampling.
@@
Both the above approaches treat the one-step backward update as independent across dimensions, which accumulates simulation errors over the course of the backward process. To combat this problem, additional corrector steps can be used to ensure the marginal distribution of samples $x_{t}$ at time $t$ matches $q_{t}$. Campbell et al. (2022) show that a corrector process with rates,
@@
leaves the stationary distribution $q_{t}$ invariant. In practice, one must use the learned backward rates $\tilde{R}_{t}^{\theta}$ as an approximation for the true backward rates $\tilde{R}_{t}$. This can be thought of as an “uninformed” corrector, as it is simply a summation of the forward and backward rates, neither of which explicitly targets problematic tokens in the masked diffusion setting. Nevertheless, Campbell et al. (2022) show that this uninformed corrector increases sample quality in practice.
@@ 3.1 Informed Correctors
To design a corrector scheme for our sampler, we first identify a Markov chain that has the marginal distribution $q_{t}$ as a stationary distribution. For discrete distributions, a straightforward option is to use Gibbs sampling, [^1] Other Markov chains could be employed instead. For example, the forward-backward corrector proposed by Campbell et al. (2022) leverages a birth-death process that has $q_{t}$ as the stationary distribution (Shi et al., 2022). We also tested locally-balanced Markov processes (Zanella, 2019), but their performance was inferior to Gibbs. i.e., we iteratively select an index from $d=1,\ldots,D$ and resample the $d$-th dimension according to the conditional distribution
@@
To improve efficiency, we use an adaptive random-scan Gibbs sampler with a non-uniform selection distribution (Levine and Casella, 2006; Levine et al., 2005; Łatuszyński et al., 2013; Mitliagkas and Mackey, 2017). We choose this distribution so that the corrector prioritizes updating those dimensions that are more likely to have simulation errors. To achieve this, we let $c_{d}$ be a confidence score indicating the likelihood of correctness in the $d$-th dimension of the current sample. We select the indices to update in the next $k$ steps $d_{1},\dots,d_{k}$ by sampling without replacement from the distribution $\mathrm{Cat}(\frac{\exp(-c_{d}/\tau)}{\sum_{d^{\prime}}\exp(-c_{d^{\prime}}/\tau)})$, where $\tau$ is a temperature parameter. This sample can be generated using the Gumbel-top-$k$ trick (Maddison et al., 2014; Kool et al., 2019): First draw $D$ independent Gumbel random variates, $g_{d}\sim\mathrm{Gumbel}(0,1)$, one for each dimension. Then rank the dimensions in descending order of $c_{d}/\tau+g_{d}$ and select the top $k$ dimensions.
@@
Similar to in (Kim et al., 2025), the advantage of this definition is that it maintains low confidence for a dimension when there is an alternative token that the model assigns comparable high probabilities (i.e., $\max_{x^{\prime d}\neq x^{d}}\log q_{t}(x^{\prime d}\mid x^{\setminus d})$ is equally large). In practice, we often use this definition of the confidence as we found it consistently outperforms the former.
@@
Rather than updating the $k$ dimensions sequentially, we update them in parallel, as in Hogwild Gibbs sampling (Johnson et al., 2013). In practice, we want to keep $k$ small to avoid errors in the parallel resampling, while keeping it large enough so that errors can be fixed reliably. We study the effects of this hyperparameter in section˜5.3.
@@ 3.2 Parameterizing the Conditional Distribution Using Hollow Transformers
We solve this problem with a hollow transformer (Sun et al., 2022), a two-stream self-attention architecture where each stream runs causal attention in a different direction. The embeddings are then offset by one position so that the embedding at position $d$ does not attend to the corresponding input position. The two streams are then summed together and sent through multilayer perceptron layers to the output. By construction, the hollow transformer satisfies the property that for all $(t,\theta)$,
@@ 3.3 Learning the Hollow Transformer with the ELBO
We include a full derivation of this result in appendix˜A and show that it is related to $\mathcal{L}_{\mathcal{M}}$ via a simple label-switching trick. Notably, Campbell et al. (2022) and Shi et al. (2024) both avoid writing the loss in this form, since evaluating $p^{\theta}_{0\mid t}(x^{d}\mid M^{d}(x))$ for each different $d$ requires a separate call to the denoising network, which is prohibitively inefficient in practice. The hollow transformer fixes this issue, as now the inner sum over $d$ can be evaluated with one forward pass of the network.
@@ 4 Related Work
Alongside the development of discrete diffusion models, other classes of non-autoregressive models employing similar iterative generation principles have emerged (Savinov et al., 2021; Chang et al., 2022; Sun et al., 2022). Among these, MaskGIT (Chang et al., 2022) is notably similar to masked diffusion models (Shi et al., 2024; Zheng et al., 2025), particularly as both are trained using denoising cross-entropy losses. However, key differences distinguish them: MaskGIT utilizes a non-likelihood-based weighting for its loss function and employs a confidence-based decoding order. In contrast, masked diffusion models typically unmask tokens in a random order that aligns with their theoretical backward process.
@@
Campbell et al. (2022) and Gat et al. (2024) introduced the forward-backward and DFM correctors for discrete diffusion. However, these “uninformed” correctors can only address errors through inefficient random re-masking. ReMDM (Wang et al., 2025) proposed the remasking diffusion process that includes forward-backward and DFM as special cases, while effectively combining the predictor and corrector steps into one. Relatedly, DPC (Lezama et al., 2022a) proposed a corrector scheme for MaskGIT sampling. Although their method can be seen as “informed”, they reportedly require hundreds of steps to achieve good sample quality on ImageNet $\times256$.
@@ 5.1 Hidden Markov Modeling
and tries to predict the conditional denoising likelihood $q_{0\mid t}(x_{0}^{d}\mid x^{\setminus d}_{t})$. This can be seen as performing inference in a hidden Markov model (HMM), where the prior is the Markov chain $q_{0}$, and the observation model is $q_{t\mid 0}$. Given this, we can perform standard message-passing inference (Murphy, 2012) to find the posterior $q_{0\mid t}(x_{0}^{d}\mid x^{\setminus d}_{t})$.
@@
In this illustrative experiment we use vocabulary size $S=4$ and sequence length $D=128$ and set the “stickiness” $p$ to $0.9$. The masking schedule $\alpha_{t}$ is set to a cosine schedule as in Shi et al. (2024). Hyperparameters for informed correctors are tuned for each different number of function evaluations (NFE) setting (see appendix˜D). For samplers with correctors, we take one corrector step after each predictor step. The NFE is then equal to the sum of the number of predictor steps and corrector steps.
We show the mean and standard deviation of error rates achieved by different sampling strategies over 5 random seeds in Fig. 2. We find that ancestral sampling with the informed corrector yields the least error for all examined NFE settings. Meanwhile, the uninformed corrector consistently performs worse than using no corrector at all. This shows that an uninformed corrector step is not as valuable as an additional predictor step in this setting. We demonstrate this pictorially in fig.˜2.
@@
MaskGIT (Lezama et al., 2022b) 227M 18 6.56 203.6 0.79 0.48
DPC (Lezama et al., 2022a) 391M 180 4.45 244.8 0.78 0.52
VQ-Diffusion (Gu et al., 2022) 518M 100 11.89 - - -
ReMDM (Wang et al., 2025) 227M 16 7.40 145.27 - -

### v5

@@ 1 Introduction
Denoising diffusion models are powerful generative models for high-dimensional data [1, 2, 3]. The central idea of diffusion models is to gradually corrupt data into noise using a “noising” or “forward” process and then to train a parameterized model (usually a deep neural network) to learn its time reversal, commonly known as the “denoising” or “backward” process. In continuous domains like image generation, the forward process is usually defined by gradual injection of Gaussian noise and scaling that transform the data distribution into a standard normal. The backward process is then approximated by learning the gradient of the log density (also known as the “score function”) of the marginal distribution. To draw samples from a trained model, one first generates Gaussian noise and then simulates the backward process using the score information. Diffusion generative models are currently the dominant framework for image generation, and they can generate high-resolution images with stunning details given prompts from a user [4, 5].
@@
Given the success of diffusion models for continuous data, recent work has explored diffusion modeling in discrete domains [6, 7, 8, 9, 10, 11, 12]. Discrete diffusion models can be applied to language [11], protein sequences [13, 14], graphs [15], and more. Notably, Campbell et al. [8] developed a general framework for discrete denoising diffusion in continuous time. They formulated the forward and backward processes as continuous time Markov chains (CTMCs), and they learned to predict the distribution over denoised data via an evidence lower bound objective. Concurrent work by Shi et al. [12], Sahoo et al. [16] and Ou et al. [17] focused on discrete diffusion with the absorbing state forward process (a.k.a., masked discrete diffusion), deriving a simple formulation of the ELBO that reduces to a weighted cross-entropy. Shi et al. [12] also noted that this simple objective provides a unifying perspective to previous works [8, 11, 18], all of which are parameterizations and modifications of the same ELBO objective.
@@
Despite conceptual advances in discrete diffusion, the efficiency-accuracy trade-off in simulating the continuous-time backward process still limit their effectiveness. Song et al. [19] proposed using approximate MCMC correctors to fix the discretization error in simulating the backward process. Campbell et al. [8] extended the predictor-corrector sampling to discrete diffusion models and discovered that the corrector steps significantly impact generative performance. In practice however, a clear understanding of how corrector steps contribute to sample quality is still missing, and recent works [11, 12, 17, 16] still use predictor-only samplers for generation.
@@
We show that predictor-corrector schemes can be improved by leveraging different sampling methods. Analyzing the forward-backward corrector used by Campbell et al. [8], we identify its failure mode in masked diffusion and illustrate it with a simple Markov chain modeling task. To address this shortcoming, we propose the informed corrector — an alternative corrector scheme inspired by adaptive Gibbs sampling that fixes the issues of the forward-backward corrector and explicitly targets low-probability dimensions. Sampling with the informed corrector drastically decreases error rate in our synthetic example, as well as in Text8, where the model makes much fewer spelling errors than baselines. On tokenized ImageNet $256\times 256$, we show that the informed corrector yield diverse, high quality samples competitive with existing diffusion models.
@@ 2.1 Continuous-Time Discrete Diffusion Models
Consider a data vector $x_{0}\in\mathcal{S}^{D}$ sampled from a data distribution $p_{\text{data}}$, where $\mathcal{S}$ is the base space for each component. Denoising diffusion models [1, 2, 3, 19] model this unknown data distribution by progressively adding noise to data samples until their distribution converges to a simple target, such as a uniform distribution. Then, a deep neural network is trained to invert this process and remove the noise. This reverse process forms a generative model from which new samples can be drawn. The noising and generative processes are also called the forward and backward processes, respectively. We will adopt this convention for the rest of the paper.
@@
Campbell et al. [8] approximated the backward rate by learning a parameterized denoising model $p^{\theta}_{0|t}(x_{0}\mid y)\approx q_{0|t}(x_{0}\mid y)$. Alternatively, Meng et al. [20] and Lou et al. [11] observed that the ratio $s_{t}(y)_{x}\triangleq\nicefrac{{q_{t}(x)}}{{q_{t}(y)}}$ plays a similar role as the continuous score function $\nabla\log q_{t}(y)$, up to an additive constant. Therefore, they proposed to learn an approximate score function instead, $s^{\theta}_{t}(y)$.
@@
for a base rate matrix $R_{b}\in\mathbb{R}^{S\times S}$, absorbing the time dependence into the scalar coefficient $\beta(t)$. Here, $x^{\setminus d}$ represents all components of $x$ apart from the $d$-th dimension. Backward rates are also factorized with this forward rate factorization, allowing the objective in ˜2 to be nicely decomposed into a sum over dimensions, as detailed by Campbell et al. [8].
@@ 2.2 Absorbing Discrete Diffusion
For generic discrete data, two forward processes are most commonly used: the uniform process and the absorbing state process [7]. We focus on the absorbing state process in this paper, for which the base rate matrix for each dimension can be written as
@@
where we augment the vocabulary set $\mathcal{S}$ by introducing the Mask token, which represents the absorbing state. The absorbing state diffusion is important to consider because it is very commonly used to model sequence data without ordinal structure (e.g., text and protein sequences). It is also intimately connected to masked language models [7, 21, 22].
@@
Recently, Shi et al. [12] introduced the MD4 diffusion model, which outperformed previous discrete diffusion models on both text and pixel-level image generation benchmarks. Furthermore, they noted that several continuous time denoising diffusion objectives (including ˜2 ) are mathematically equivalent and can be greatly simplified in the case of masked diffusion. The simplified MD4 loss is:
@@ 2.3 Sampling From the Backward Process
Campbell et al. [8] proposed to use tau-leaping, an approximate sampling method that allows simultaneous updates to multiple dimensions for a single model evaluation. Tau-leaping assumes that the backward process transition rate is constant between $t-\tau,t)$, regardless of whether states change during that interval. Under this assumption, multiple transition events in the interval can be applied simultaneously. Leverging the structure of masked diffusion, MD4 [12] used a simpler approach. Instead of trying to simulate the continuous time process, they directly sampled from the finite-time backward conditional distribution $p^{\theta}_{t-\Delta t\mid t}(\,\cdot\mid x_{t})$. They call this approach ancestral sampling.
@@
Both the above approaches treat the one-step backward update as independent across dimensions, which accumulates simulation errors over the course of the backward process. To combat this problem, additional corrector steps can be used to ensure the marginal distribution of samples $x_{t}$ at time $t$ matches $q_{t}$. Campbell et al. [8] show that a corrector process with rates,
@@
leaves the stationary distribution $q_{t}$ invariant. In practice, one must use the learned backward rates $\tilde{R}_{t}^{\theta}$ as an approximation for the true backward rates $\tilde{R}_{t}$. This can be thought of as an “uninformed” corrector, as it is simply a summation of the forward and backward rates, neither of which explicitly targets problematic tokens in the masked diffusion setting. Nevertheless, Campbell et al. [8] show that this uninformed corrector increases sample quality in practice.
@@ 3.1 Informed Correctors
To design a corrector scheme for our sampler, we first identify a Markov chain that has the marginal distribution $q_{t}$ as a stationary distribution. For discrete distributions, a straightforward option is to use Gibbs sampling, [^1] Other Markov chains could be employed instead. For example, the forward-backward corrector proposed by Campbell et al. [8] leverages a birth-death process that has $q_{t}$ as the stationary distribution [23]. We also tested locally-balanced Markov processes [24], but their performance was inferior to Gibbs. i.e., we iteratively select an index from $d=1,\ldots,D$ and resample the $d$-th dimension according to the conditional distribution
@@
To improve efficiency, we use an adaptive random-scan Gibbs sampler with a non-uniform selection distribution [25, 26, 27, 28]. We choose this distribution so that the corrector prioritizes updating those dimensions that are more likely to have simulation errors. To achieve this, we let $c_{d}$ be a confidence score indicating the likelihood of correctness in the $d$-th dimension of the current sample. We select the indices to update in the next $k$ steps $d_{1},\dots,d_{k}$ by sampling without replacement from the distribution $\mathrm{Cat}(\frac{\exp(-c_{d}/\tau)}{\sum_{d^{\prime}}\exp(-c_{d^{\prime}}/\tau)})$, where $\tau$ is a temperature parameter. This sample can be generated using the Gumbel-top-$k$ trick [29, 30]: First draw $D$ independent Gumbel random variates, $g_{d}\sim\mathrm{Gumbel}(0,1)$, one for each dimension. Then rank the dimensions in descending order of $c_{d}/\tau+g_{d}$ and select the top $k$ dimensions.
@@
Similar to in [31], the advantage of this definition is that it maintains low confidence for a dimension when there is an alternative token that the model assigns comparable high probabilities (i.e., $\max_{x^{\prime d}\neq x^{d}}\log q_{t}(x^{\prime d}\mid x^{\setminus d})$ is equally large). In practice, we often use this definition of the confidence as we found it consistently outperforms the former (see table˜4).
@@
Rather than updating the $k$ dimensions sequentially, we update them in parallel, as in Hogwild Gibbs sampling [32]. In practice, we want to keep $k$ small to avoid errors in the parallel resampling, while keeping it large enough so that errors can be fixed reliably. We study the effects of this hyperparameter in section˜5.3.
@@ 3.2 Parameterizing the Conditional Distribution Using Hollow Transformers
We solve this problem with a hollow transformer [33], a two-stream self-attention architecture where each stream runs causal attention in a different direction. The embeddings are then offset by one position so that the embedding at position $d$ does not attend to the corresponding input position. The two streams are then summed together and sent through multilayer perceptron layers to the output. By construction, the hollow transformer satisfies the property that for all $(t,\theta)$,
@@ 3.3 Learning the Hollow Transformer with the ELBO
We include a full derivation of this result in appendix˜A and show that it is related to $\mathcal{L}_{\mathcal{M}}$ via a simple label-switching trick. Notably, Campbell et al. [8] and Shi et al. [12] both avoid writing the loss in this form, since evaluating $p^{\theta}_{0\mid t}(x^{d}\mid M^{d}(x))$ for each different $d$ requires a separate call to the denoising network, which is prohibitively inefficient in practice. The hollow transformer fixes this issue, as now the inner sum over $d$ can be evaluated with one forward pass of the network.
@@ 4 Related Work
Alongside the development of discrete diffusion models, other classes of non-autoregressive models employing similar iterative generation principles have emerged [34, 35, 33]. Among these, MaskGIT [35] is notably similar to masked diffusion models [12, 36], particularly as both are trained using denoising cross-entropy losses. However, key differences distinguish them: MaskGIT utilizes a non-likelihood-based weighting for its loss function and employs a confidence-based decoding order. In contrast, masked diffusion models typically unmask tokens in a random order that aligns with their theoretical backward process. Inspired by MaskGIT, large language diffusion models such as LLaDA [37] adopt similar heuristics for remasking tokens at each prediction step. This effectively alters the predictor’s unmasking order based on confidence, which can introduce bias into the sampling process, causing it to deviate from the diffusion model’s theoretical reverse process. In contrast, our confidence metric is only used to define the selection distribution of a random-scan Gibbs sampler, and the informed corrector preserves the marginal distributions of the original diffusion framework to first order.
@@
Campbell et al. [8] and Gat et al. [38] introduced the forward-backward and DFM correctors for discrete diffusion. However, these “uninformed” correctors can only address errors through inefficient random re-masking. ReMDM [39] proposed the remasking diffusion process that includes forward-backward and DFM as special cases, while effectively combining the predictor and corrector steps into one. Relatedly, DPC [40] proposed a corrector scheme for MaskGIT sampling. Although their method can be seen as “informed”, they reportedly require hundreds of steps to achieve good sample quality on ImageNet $\times256$.
@@ 5.1 Hidden Markov Modeling
and tries to predict the conditional denoising likelihood $q_{0\mid t}(x_{0}^{d}\mid x^{\setminus d}_{t})$. This can be seen as performing inference in a hidden Markov model (HMM), where the prior is the Markov chain $q_{0}$, and the observation model is $q_{t\mid 0}$. Given this, we can perform standard message-passing inference [41] to find the posterior $q_{0\mid t}(x_{0}^{d}\mid x^{\setminus d}_{t})$.
@@
We show error rates (mean and standard deviation over 5 random seeds) in Fig. 2. We find that ancestral sampling with the informed corrector yields the least error for all examined NFE settings. Meanwhile, the uninformed corrector consistently performs worse than using no corrector at all, showing that an uninformed corrector step is not as valuable as an additional predictor step in this setting (see fig.˜2). More experimental details can be found in appendix˜D.
@@
MaskGIT [42] 227M 18 6.56 203.6 0.79 0.48
DPC [40] 391M 180 4.45 244.8 0.78 0.52
VQ-Diffusion [43] 518M 100 11.89 - - -
ReMDM [39] 227M 16 7.40 145.27 - -